Calculate the mass of an atom of molecules of carbon dioxide
 

Grade:11

3 Answers

Arun
4 years ago
Mass of 1 atom of carbon is 12 a.u.
Mass of 2 atoms of oxygen is 2*16 a.u. = 32 a.u.
Hence the mass of one molecule of CO2 = 32+12= 44 a.u.
Vikas TU
3 years ago
We can calculate the mass of one molecule of carbon dioxide by adding together the masses of 1 atom of carbon and 2 atoms of oxygen. The molecular mass of a compound is the mass of one molecule of that compound. The molecular mass of carbon dioxide is 44.01amu
Yashraj
3 years ago
Mass of 1 atom of carbon = 12 a.u.
 Mass oxygen is 2×16 a.u. = 32 a.u
 mass of one molecule of CO2 = 
44 a.u.